# Lens: A Foundation Model for Network Traffic

### v1

@@ Abstract
Network traffic refers to the amount of information being sent and received over the internet or any system that connects computers. Analyzing and understanding network traffic is vital for improving network security and management. However, the analysis of network traffic poses great challenges due to the unique characteristics of data packets, such as heterogeneous headers and encrypted payload lacking semantics. To capture the latent semantics of traffic, a few studies have adopted pre-training techniques based on the Transformer encoder or decoder to learn the representations from large-scale traffic data. However, these methods typically excel only in traffic understanding (classification) or traffic generation tasks. To address this issue, we develop Lens, a foundational network traffic model that leverages the T5 architecture to learn the pre-trained representations from large-scale unlabeled data. Harnessing the strength of the encoder-decoder framework, which captures the global information while preserving the generative ability, our model can better learn the representations from large-scale network traffic. To further enhance pre-training performance, we design a novel loss that integrates three distinct tasks, namely Masked Span Prediction (MSP), Packet Order Prediction (POP), and Homologous Traffic Prediction (HTP). Evaluation results on multiple benchmark datasets demonstrate that the proposed Lens outperforms the baselines in most downstream tasks related to both traffic understanding and traffic generation. Notably, it also requires considerably less labeled data for fine-tuning compared to current methods.
@@ 1 Introduction
Over the past decades, many approaches have been developed for network traffic analysis. Early works [36, 1, 28, 42] mainly utilized statistical methods, heavily dependent on manually crafted features. To address this issue, some studies have employed deep learning methods [34, 21, 23, 39] to extract complicated features from raw data. While these approaches have shown impressive results in specific tasks, they often require extensive labeled data and struggle with generalization to novel tasks. Consequently, few recent works have explored pre-training techniques to learn the latent representations from large-scale unlabeled traffic data. For instance, some works have introduced BERT-based models to learn the pre-trained representations, subsequently applied to traffic understanding (classification) in downstream tasks. Despite their effectiveness in traffic classification, these models lack generative capabilities due to their encoder-only architecture. Indeed, GPT-based models [29, 30, 3] excel in generation capability, but they only learn from previous information in a sequence without considering the following information. As a result, this unidirectional learning technique may not capture the global information of input data, impeding their capacity to fully understand network traffic. Thus, a pivotal question arises: how can we achieve good performance of both traffic understanding and generation within the scope of network traffic analysis?
@@
Figure 1: The overall framework of Lens. It consists of three parts: traffic tokenization, pre-training using the T5 Architecture, and fine tuning.
@@
In this paper, we develop Lens, a foundation model that leverages the T5 architecture [31, 41] to learn the pre-trained representations from massive traffic data. Thanks to the encoder-decoder architecture of T5, the proposed model excels in both traffic understanding and traffic generalization. As illustrated in Fig. 1, the overall framework consists of three stages: tokenization, pre-training, and fine-tuning. For traffic tokenization, considering the heterogeneous headers and payloads that may contain plaintext and encrypted data, we standardize different traffic types by converting their content into hexadecimal format. Furthermore, we explore the impact of different tokenization methods, such as SentencePiece and WordPiece, on the model performance. At the pre-training stage, we design a new loss function that addresses three tasks, including Masked Span Prediction (MSP), Packet Order Prediction (POP), and Homologous Traffic Prediction (HTP), to better discern the latent patterns from various network traffic. Specifically, MSP serves as a denoising task that prompts Lens to recover masked tokens, thereby enabling the model to capture the representations within network traffic. POP predicts the order of packets within a single traffic input, facilitating the capture of latent features within the same session flow, where flow means a sequence of packets exchanged between two endpoints. Inspired by [19], HTP aids in identifying the relational patterns across different flows. Regarding fine-tuning, we leverage task-oriented prompts coupled with a small amount of task-specific labeled data to delicately tune the parameters of the proposed Lens, making it effectively adaptive to various downstream tasks.
@@
We evaluate the performance of Lens on both understanding and generation tasks across 6 real-world datasets [9, 11, 40, 37, 25, 27], including 15 traffic understanding tasks and 5 header fields generation tasks. Compared to ET-BERT, we use 50% $\sim$ 95% less labeled data for fine-tuning. Extensive experimental results demonstrate that our model can reach 10.75% and 0.59% higher accuracy than the baselines for packet-level and flow-level traffic understanding, respectively. For traffic generation, the proposed Lens has lower Jensen–Shannon Divergence (JSD) and Total Variation Distance (TVD) than baselines, which can be reduced by $0.1353$ and $0.0825$ on average.
@@
In summary, our contributions are four fold: (1) we introduce Lens, the first large network traffic model using the T5 architecture for both traffic understanding and generation; (2) we investigate the impact of different tokenization methods and their settings on the pre-trained representations of network traffic; (3) we integrate three distinct tasks into the model’s pre-training stage to learn better representations from raw traffic data; (4) Extensive experiments show the superiority of our model over baselines in various downstream tasks related to traffic understanding and generation.
@@ 2.1 Network Traffic Understanding
Pre-training Approaches. To improve model generalization ability, recent studies have adopted pre-training techniques to learn the representations from large-scale traffic data in an unsupervised manner. For instance, PERT [12] and ET-BERT [19] leveraged the pre-trained ALBERT [17] and BERT [8] models to learn the latent representations, respectively. However, these models are not applicable to traffic generation due to their encoder-only structure. Most recently, NetGPT [24] built on GPT-2 [30] has been developed for both traffic classification and generation in a generative way. However, it fails to capture the global information of input data due to its decoder-only framework, leading to sub-optimal performance on traffic understanding tasks [41]. Moreover, NetGPT [^1] NetGPT is currently preprinted and not open-sourced, we thus do not compare our method with it. pre-trained on network traffic without removing IP addresses, which may result in a risk of privacy leakage. Lastly, it is only implemented on 8 downstream tasks for traffic classification while our Lens is tested on 15 downstream tasks. Table 1 summarizes the comparison of the proposed Lens  and other existing methods.
Method Encoder Decoder Curated PT Fun IP Masking GT
@@ 2.2 Network Traffic Generation
Pre-training-based Generation. Most recently, researchers employed pre-training technique for traffic generation. Only one work called NetGPT has been developed to generate key header fields. However, NetGPT did not compare it with state-of-the-art methods like NetShare in the experiments. Thus, it is hard to assess its performance on generalization.
@@ 3 Proposed Lens
In this section, we will introduce Lens, a large network traffic model for both traffic understanding and generation. Fig. 1 illustrates the overall framework of the proposed model, consisting of three main stages: traffic tokenization, pre-training, and fine-tuning. In Section 3.1, we will describe three different tokenization approaches: Vanilla Vocab, SentencePiece [15], and WordPiece [8]. We try to figure out which one performs best in our experiments in Sec. 4. Then we will present the pre-training process that involves three tasks: Masked Span Prediction (MSP), Packet Order Prediction (POP), and Homologous Traffic Prediction (HTP) in Section 3.2. Lastly, Section 3.3 will focus on the fine-tuning of the pretrained Lens on a small amount of labeled data.
@@ 3.1 Traffic Tokenization
3.1.1 Data pre-processing.
3.1.2 Tokenization.
As shown in Table 2, we further explore different settings for SentencePiece and WordPiece. Specifically, for SentencePiece, we examine two configurations: ‘SentencePiece String’, which processes an entire hex flow as a single string, and ’SentencePiece Word’, which first divides hex flows into 4-digit words before treating it as a string. For WordPiece, the focus is on ‘WordPiece Word’, where the hex traffic is split into 4-digit words without using a predefined vocabulary while ‘WordPiece Pd.’ uses ‘Vanilla Vocab’ as the predefined vocabulary.
@@ 3.2.2 New Objective with Three Tasks
where $\theta$ represents model parameters, $k$ is the total number of masked spans, $\text{MSAPN}_{i}$ is $i$-th span generated by the decoder part of Lens, $\text{SPAN}_{i}$ denotes the original span tokens, $\mathbf{x}_{\text{in}}$ is the input sequence of the encoder, $\text{MSAPN}_{<i}$ indicates the generated spans from the decoder before $i$. $\text{MSAPN}_{i}$ is generated given $\mathbf{x}_{\text{in}}$ and $\text{MSAPN}_{<i}$ when parameterized by $\theta$.
@@
Given an input flow $f_{i}$ (the $i$-th sample), suppose that it is composed of $t$ packets $\sum^{t}P^{(i)}_{k}(t<=3)$, where $P_{k}^{(i)}$ denotes the $k$-th packet from flow $i$. Let $p_{ij}$ represent the softmax probability model output of the $j$-th <pkt> token and $y_{ij}$ denote whether $j$-th packet is on the $j$-th position of the original flow, where $y_{ij}\in\{0,1\}$. Besides, we use $z_{i}\in\{0,1\}$ to denote whether packets from $f_{i}$ come from homologous traffic (which will be introduced in the next paragraph). This can help us avoid applying both POP and HTP tasks on the same flow since applying POP to the heterologous flow is meaningless. Then we have the loss function for POP as follows
$\mathcal{L}_{\text{POP}}=-\sum_{i=1}^{M}z_{i}\sum_{j=1}^{t}y_{ij}\log p_{ij}$ (2)
@@ 3.3 Fine-Tuning
After finishing model pre-training, the next step is to fine tune the proposed Lens on a small amount of data from various downstream tasks. We can achieve this by using task-specific transfer learning or multi-task learning. In the following, we detail how to fine tune our foundation model in traffic understanding and traffic generation, respectively.
@@ 4.1 Experimental Settings
At the beginning of our experiments, we first split raw pcap files into individual flows, which are then converted into a hexadecimal format. Then, we randomly choose some flows from each dataset for fine tuning, as detailed in Table 3. The remaining of data is used for pre-training. Below, we introduce the detailed experimental setup for pre-training and fine tuning.
@@
In contrast to pre-training that only use flow-level data, fine tuning may input a flow or a packet into the model since some downstream tasks operate at either a flow level or packet level. Specifically, for flow-level tasks, we extract the first three packets from each large flow as mentioned earlier. For packet-level tasks, we choose the first five packets of each flow. This selective strategy demonstrates the Lens’s adaptability and effective performance even with limited labeled data. Note that compared to other methods like ET-BERT, our model requires 50% $\sim$ 95% less packets for fine tuning. Regarding traffic generation, we mask the header fields that would be predicted during both training and testing. At the fine tuning stage, we split pre-processed data into a training set and a testing set at a ratio of 4:1.
@@ 4.1.2 Evaluation Metrics
FS-Net 0.1563 0.1046 0.1068 0.0574 0.4440 0.2050 0.5980 0.3742 0.8800 0.8714 0.8375 0.4558 0.4883 0.4250 0.4908
BiLSTM_Att 0.6083 0.4428 0.1078 0.1006 0.6049 0.5415 1.0000 1.0000 0.6506 0.6400 0.9001 0.7652 0.7570 0.5943 0.6143
Datanet 0.5743 0.3278 0.0079 0.0001 0.5494 0.4350 1.0000 1.0000 0.6558 0.6403 0.8934 0.6147 0.7633 0.5863 0.5796
DeepPacket 0.5779 0.3630 0.0579 0.0473 0.5553 0.4666 1.0000 1.0000 0.6253 0.5862 0.9015 0.7673 0.7500 0.5736 0.5984
TSCRNN 0.6118 0.4368 0.1043 0.0964 0.5935 0.5204 1.0000 1.0000 0.6502 0.6443 0.8968 0.6373 0.7578 0.6067 0.6185
YaTC 0.9890 0.9866 0.6248 0.5938 0.9503 0.9263 1.0000 1.0000 0.9800 0.9799 0.9647 0.9265 0.8394 0.7673 0.8540
ET-Bert (flow) 0.7454 0.7136 0.3401 0.2992 0.8650 0.8597 1.0000 1.0000 0.6714 0.6279 0.8384 0.7035 0.7026 0.5662 0.6984
ET-Bert (pkt) 0.9845 0.9824 0.9289 0.9254 0.9872 0.9887 1.0000 1.0000 1.0000 1.0000 0.9704 0.9480 0.9320 0.9181 0.9697
Lens(flow) 0.9822 0.9808 0.9189 0.9143 0.9867 0.9816 1.0000 1.0000 0.9506 0.9507 0.9712 0.9285 0.9054 0.9061 0.9615
Lens(pkt) 0.9832 0.9806 0.9349 0.9266 0.9896 0.9880 0.9982 0.9981 0.9885 0.9885 0.9841 0.9658 0.9412 0.9363 0.9756
Table 5: Comparison results of traffic understanding tasks from task 9 to task 15.
4.1.3 Implementation Details.
We adopt Google’s T5-v1.1-base model [31], which contains roughly 0.25B parameters, as the foundational architecture for the proposed Lens. For the final objective function introduced in Eq. 4, we set both $\alpha$ and $\beta$ to $0.2$ based on grid search method. At the pre-training stage, we set the batch size to $128$ and implement gradient accumulation over $4$ training steps. The learning rate is set to $2\times 10^{-2}$ over a total of 56,000 steps with a warmup steps of 8000. We also use a dropout rate of $0.1$ to mitigate the risk of overfitting. For fine-tuning, we set the batch size to 32 and the learning rate to $3\times 10^{-5}$, training each downstream task for 10 epochs using the AdamW optimizer [22] on a single GPU with 48GB memory capacity.
4.1.4 Downstream tasks.
To evaluate the good performance of our Lens, we select 15 different understanding tasks to classify network traffic as well as 5 generation tasks for generating important header fields (source IP, destination IP, source port, destination port, packet length). To our best knowledge, we evaluate a more wide range of downstream tasks compared to existing works. Below, we describe the experimental settings for both traffic understanding and generation.
Traffic Generation Tasks: Based on the methodology of NetShare [44], we generate five vital header fields at the packet-level as follows: source IP, destination IP, source port, destination port, and packet length. The generation of these synthetic header fields can help create usable pcap traffic for network simulation. For each dataset in our experiments, all generation tasks are individually conducted using a supervised fine-tuning approach.
4.2 Evaluation Results
In this subsection, we first assess the proposed Lenson traffic understanding tasks, and then apply it to traffic generation tasks. For traffic understanding, we compare the proposed model against 7 state-of-the-art (SOTA) baselines. For traffic generation, we primarily compare our method with the most competitive and open-sourced baseline NetShare.
@@ 4.2 Evaluation Results
First, we evaluate the performance of Lens on a variety of traffic understanding tasks. We compare it with the following baselines: FS-Net [21], BiLSTM_Att [43], Datanet [39], DeepPacket [23], YaTC [45], TSCRNN [18], and ET-Bert [19]. In this experiment, we assess different methods with a focus on two levels of data granularity: flow-level and packet-level. Tables 4 and 5 illustrate the comparison results of our Lens and the baselines in traffic understanding tasks. From the evaluation results, we can conclude with three main insights: (1) Lensconsistently achieves SOTA performance in most of tasks, indicating its superiority overall the baselines. (2) With limited testing samples, Lens using packet-level input reaches higher accuracy and F1 than that of using flow-level input. In particular, ET-BERT demonstrates a significant average accuracy difference of 30.13% between these two data granularities. In contraset, our Lens still achieves comparable performance, with only 1.41% variance across two levels of data granularity. (3) Even if without CrossPlatform dataset in pre-training, Lens outperforms the baselines in three out of four tasks, especially for tasks 8, 10, and 11, demonstrating its strong generalization ability.
@@
In addition, we conduct experiments to evaluate the performance of Lens on traffic generation tasks by comparing it with the SOTA method, NetShare [44]. In this experiment, two evaluation metrics: Jensen-Shannon Divergence (JSD) and Total Variation Distance (TVD), are used to measure the results. In terms of metrics JSD and TVD, a lower value means better performance, indicating smaller distribution difference between the generated and ground-truth data. The generation results related to port number and packet length are illustrated in Tables 6, while the results of IP generation are introduced in Appendix A.1. We draw some conclusions from this table as follows: (1) Lensoutperforms NetShare in most tasks in terms of TVD and JSD, demonstrating its good generation performance. Compared to NetShare, Lenscan generate more high-quality synthetic traffic data by reducing JSD and TVD by 62% and 77% on average, respectively. (2) Lensconsistently outperforms the NetShare in the task of packet length generation under both metrics. (3) Even if our model does not per-train on CrossPlatform dataset, it still surpasses the NetShare across most generation tasks in this dataset. More visualization results are presented in Appendix A.2.

### v3

@@ Abstract
Network traffic refers to the amount of data being sent and received over the internet or any system that connects computers. Analyzing and understanding network traffic is vital for improving network security and management. However, the analysis of network traffic is challenging due to the diverse nature of data packets, which often feature heterogeneous headers and encrypted payloads lacking semantics. To capture the latent semantics of traffic, a few studies have adopted pre-training techniques based on the Transformer encoder or decoder to learn the representations from massive traffic data. However, these methods typically excel in traffic understanding (classification) or traffic generation tasks. To address this issue, we develop Lens, a foundation model for network traffic that leverages the T5 architecture to learn the pre-trained representations from large-scale unlabeled data. Harnessing the strength of the encoder-decoder framework, which captures the global information while preserving the generative ability, our model can better learn the representations from raw data. To further enhance pre-training effectiveness, we design a novel loss that combines three distinct tasks: Masked Span Prediction (MSP), Packet Order Prediction (POP), and Homologous Traffic Prediction (HTP). Evaluation results across various benchmark datasets demonstrate that the proposed Lens outperforms the baselines in most downstream tasks related to both traffic understanding and generation. Notably, it also requires much less labeled data for fine-tuning compared to current methods.
@@ 1 Introduction
Over the past decades, many approaches have been developed for network traffic analysis. Early works [36, 1, 28, 42] mainly utilized statistical methods, heavily dependent on manually crafted features. To address this issue, some studies have employed deep learning methods [34, 21, 23, 39] to extract complicated features from raw data. While these approaches have shown impressive results in specific tasks, they often require extensive labeled data and struggle with generalization to novel tasks. Consequently, few recent works have explored pre-training techniques to learn the latent representations from large-scale unlabeled traffic data. For instance, some works have introduced BERT-based models [8, 19] to learn the pre-trained representations, subsequently applied to traffic understanding (classification) in downstream tasks. Despite their effectiveness in traffic classification, these models lack generative capabilities due to their encoder-only architecture. Indeed, GPT-based models [29, 30, 3] excel in generation capability, but they only learn from previous information in a sequence without considering the following information. As a result, this unidirectional learning technique may not capture the global information of input data, impeding their capacity to fully understand network traffic. Thus, a pivotal question arises: how can we achieve good performance of both traffic understanding and generation within the scope of network traffic analysis?
@@
Figure 1: The overall framework of Lens. It consists of three parts: traffic tokenization, pre-training using the T5 Architecture, and fine-tuning.
@@
In this paper, we develop Lens, a foundation model that leverages the T5 architecture [31, 41] to learn the pre-trained representations from massive traffic data. Thanks to the encoder-decoder architecture of T5, the proposed model excels in both traffic understanding and traffic generalization. As illustrated in Fig. 1, the overall framework consists of three stages: tokenization, pre-training, and fine-tuning. During tokenization, to accommodate the diverse nature of traffic data, including both plaintext and encrypted information, we standardize different traffic types by converting their content into a hexadecimal format. Furthermore, we explore the impact of different tokenization methods, such as SentencePiece [15] and WordPiece [19], on model performance. At the pre-training stage, we design a new loss function that addresses three tasks, including Masked Span Prediction (MSP), Packet Order Prediction (POP), and Homologous Traffic Prediction (HTP), to better discern the latent patterns from various network traffic. Specifically, MSP serves as a denoising task that prompts Lens to recover masked tokens, thereby enabling the model to capture the representations within network traffic. POP predicts the order of packets within a single traffic input, facilitating the capture of latent features within the same session flow, where flow means a sequence of packets exchanged between two endpoints. Inspired by [19], HTP aids in identifying the relational patterns across different flows. For fine-tuning, we leverage task-oriented prompts coupled with a small amount of task-specific labeled data to precisely tune the parameters of Lens, making it effectively adaptive to various downstream tasks.
@@
We assess the performance of Lens in both understanding and generation tasks using 6 real-world datasets [9, 11, 40, 37, 25, 27], including 15 traffic understanding tasks and 5 header fields generation tasks. Extensive experimental results show that Lens  achieves an average accuracy improvement of 10.75% for packet level understanding tasks and 0.59% for flow level tasks, outperforming the baselines. Additionally, regarding traffic generation tasks, it outperforms the baselines in terms of Jensen–Shannon Divergence (JSD) and Total Variation Distance (TVD), reduced by an average of 0.1353 and 0.0825, respectively. Finally, our method requires 50% $\sim$ 95% less labeled data for fine-tuning compared to the state-of-the-art approach.
@@
In summary, our contributions are four-fold: (1) we introduce Lens, the first large network traffic model using the T5 architecture for traffic understanding and generation; (2) we explore the impact of different tokenization methods and their settings on the pre-training effectiveness; (3) we enhance the pre-training phase by integrating three specialized tasks, aiming to learn better representations from raw traffic data; (4) extensive experiments show the superiority of our model over baselines in various downstream tasks related to traffic understanding and generation.
@@ 2.1 Network Traffic Understanding
Pre-training Approaches. To improve model generalization ability, recent studies have adopted pre-training techniques to learn the representations from large-scale traffic data in an unsupervised manner. For instance, PERT [12] and ET-BERT [19] leveraged the pre-trained ALBERT [17] and BERT [8] models to learn the latent representations, respectively. However, these models are not applicable to traffic generation due to their encoder-only structure. Most recently, NetGPT [24] built on GPT-2 [30] has been developed for both traffic classification and generation in a generative way. However, it fails to capture the global information of input data due to its decoder-only framework, leading to sub-optimal performance on traffic understanding tasks [41]. Moreover, NetGPT [^1] NetGPT is currently preprinted and not open-sourced, and we thus do not compare our method with it. pre-trained on network traffic without removing IP addresses, which may result in a risk of privacy leakage. Lastly, it is only implemented on 8 downstream tasks for traffic classification while our Lens is tested on 15 downstream tasks. Table 1 summarizes the comparison of the proposed Lens  and other existing methods.
Method Encoder Decoder Curated PT Func IP Masking GT
@@ 2.2 Network Traffic Generation
Pre-training-based Generation. Most recently, researchers employed pre-training technique for traffic generation. Only one work called NetGPT [24] has been developed to generate key header fields. However, NetGPT did not compare it with state-of-the-art methods like NetShare in the experiments. Thus, it is hard to assess its performance on generalization.
@@ 3 Proposed Lens
In this section, we will introduce Lens, a large network traffic model for both traffic understanding and generation. Fig. 1 illustrates the overall framework of the proposed model, consisting of three main stages: traffic tokenization, pre-training, and fine-tuning. In Section 3.1, we will describe three different tokenization approaches: Vanilla Vocab, SentencePiece [15], and WordPiece [8]. We try to figure out which one performs best in our experiments in Sec. 4. Then we will present the pre-training process that involves three tasks: Masked Span Prediction (MSP), Packet Order Prediction (POP), and Homologous Traffic Prediction (HTP) in Section 3.2. Lastly, Section 3.3 will focus on the fine-tuning of the pre-trained Lens on a small amount of labeled data.
@@ 3.1 Traffic Tokenization
3.1.1 Data pre-processing.
3.1.2 Tokenization.
As shown in Table 2, we further explore different settings for SentencePiece and WordPiece. Specifically, for SentencePiece, we examine two configurations: ‘SentencePiece String’, which processes an entire hex flow as a single string, and ‘SentencePiece Word’, which first divides hex flows into 4-digit words before treating it as a string. For WordPiece, we consider two settings: (i) ‘WordPiece Word’ splits the hex traffic into 4-digit words without using a predefined vocabulary; (ii) ‘WordPiece Pd.’ uses ‘Vanilla Vocab’ as the predefined vocabulary. This predefined vocabulary provides a standardized set of words that the algorithm can use to segment and recognize patterns in the data.
@@ 3.2.2 New Objective with Three Tasks
where $\theta$ represents model parameters, $k$ is the total number of masked spans, $\text{MSPAN}_{i}$ is $i$-th span generated by the decoder part of Lens, $\text{SPAN}_{i}$ denotes the original span tokens, $\mathbf{x}_{\text{in}}$ is the input sequence of the encoder, $\text{MSPAN}_{<i}$ indicates the generated spans from the decoder before $i$. $\text{MSPAN}_{i}$ is generated given $\mathbf{x}_{\text{in}}$ and $\text{MSPAN}_{<i}$ when parameterized by $\theta$.
@@
Given an input flow $f_{i}$ (the $i$-th sample), suppose that it is composed of $t$ packets $\sum^{t}P^{(i)}_{k}(t<=3)$, where $P_{k}^{(i)}$ denotes the $k$-th packet from flow $i$. We use <pkt> token to delineate each packet. Let $y_{ij}$ be an indicator that confirms if the $j$-th packet is indeed in its original position, with 1 for yes and 0 for no, i.e, $y_{ij}=\texttt{isSamePosition}(c_{ij})$, where $c_{ij}$ is the real position for packet $j$ in flow $i$. Besides, we use $z_{i}\in\{0,1\}$ to denote whether packets from $f_{i}$ come from homologous traffic (which will be introduced in the next paragraph). This can help us avoid applying both POP and HTP tasks on the same flow since applying POP to the heterologous flow is meaningless. Then we have the following loss function for POP
$\mathcal{L}_{\text{POP}}=-\sum_{i=1}^{M}z_{i}\sum_{j=1}^{t}\log P(y_{ij}|\mathbf{x}_{\text{in}};\theta)$ (2)
@@ 3.3 Fine-Tuning
After finishing model pre-training, the next step is to fine-tune the proposed Lens on a small amount of data from various downstream tasks. We can achieve this by using task-specific transfer learning for all tasks. In the following, we detail how to fine-tune our foundation model in traffic understanding and traffic generation, respectively.
@@ 4.1 Experimental Settings
At the beginning of our experiments, we first split raw pcap files into individual flows, which are then converted into a hexadecimal format. Then, we randomly choose some flows from each dataset for fine-tuning, as detailed in Table 3. The remaining of data is used for pre-training. Below, we introduce the detailed experimental setup for pre-training and fine-tuning.
@@
In contrast to pre-training which only uses flow-level data, fine-tuning may input a flow or a packet into the model since some downstream tasks operate at either a flow level or packet level. Specifically, for flow-level tasks, we extract the first three packets from each large flow as mentioned earlier. For packet-level tasks, we choose the first five packets of each flow. This selective strategy demonstrates the Lens’s adaptability and effective performance even with limited labeled data. Note that compared to other methods like ET-BERT, our model requires 50% $\sim$ 95% fewer packets for fine-tuning. Regarding traffic generation, we mask the header fields that would be predicted during both training and testing. At the fine-tuning stage, we split pre-processed data into a training set and a testing set at a ratio of 4:1.
@@ 4.1.2 Evaluation Metrics
4.1.3 Implementation Details.
We adopt Google’s T5-v1.1-base model [31], which contains roughly 0.25B parameters, as the foundational architecture for the proposed Lens. For the final objective function introduced in Eq. 4, we set both $\alpha$ and $\beta$ to $0.2$ based on grid search method, as detailed in Appendix A.3. At the pre-training stage, we set the batch size to $128$ and implement gradient accumulation over $4$ training steps. The learning rate is set to $2\times 10^{-2}$ over a total of 56,000 steps with a warmup steps of 8000. We also use a dropout rate of $0.1$ to mitigate the risk of overfitting. For fine-tuning, we set the batch size to 32 and the learning rate to $3\times 10^{-5}$, training each downstream task for 10 epochs using the AdamW optimizer [22] on a single GPU with 48GB memory capacity.
4.1.4 Downstream tasks.
To evaluate the good performance of our Lens, we select 15 different understanding tasks to classify network traffic as well as 5 generation tasks for generating important header fields (source IP, destination IP, source port, destination port, packet length). To our best knowledge, we evaluate a wider range of downstream tasks compared to existing works. Below, we describe the experimental settings for both traffic understanding and generation.
Traffic Generation Tasks: Based on the methodology of NetShare [44], we generate five vital header fields at the packet level as follows: source IP, destination IP, source port, destination port, and packet length. The generation of these synthetic header fields can help create usable pcap traffic for network simulation. For each dataset in our experiments, all generation tasks are individually conducted using a supervised fine-tuning approach.
4.2 Evaluation Results
In this subsection, we first assess the proposed Lens on traffic understanding tasks, and then apply it to traffic generation tasks. For traffic understanding, we compare the proposed model against 7 state-of-the-art (SOTA) baselines. For traffic generation, we primarily compare our method with the most competitive and open-sourced baseline, NetShare.
@@ 4.2 Evaluation Results
Table 5: Comparison results of traffic understanding tasks from tasks 9 to 15. “Avg.” means the average accuracy on 15 tasks.
@@
FS-Net 0.1563 0.1046 0.1068 0.0574 0.4440 0.2050 0.5980 0.3742 0.8800 0.8714 0.8375 0.4558 0.4883 0.4250 0.5482
BiLSTM_Att 0.6083 0.4428 0.1078 0.1006 0.6049 0.5415 1.0000 1.0000 0.6506 0.6400 0.9001 0.7652 0.7570 0.5943 0.6321
Datanet 0.5743 0.3278 0.0079 0.0001 0.5494 0.4350 1.0000 1.0000 0.6558 0.6403 0.8934 0.6147 0.7633 0.5863 0.6123
DeepPacket 0.5779 0.3630 0.0579 0.0473 0.5553 0.4666 1.0000 1.0000 0.6253 0.5862 0.9015 0.7673 0.7500 0.5736 0.5851
TSCRNN 0.6118 0.4368 0.1043 0.0964 0.5935 0.5204 1.0000 1.0000 0.6502 0.6443 0.8968 0.6373 0.7578 0.6067 0.6212
YaTC 0.9890 0.9866 0.6248 0.5938 0.9503 0.9263 1.0000 1.0000 0.9800 0.9799 0.9647 0.9265 0.8394 0.7673 0.8655
ET-Bert (flow) 0.7454 0.7136 0.3401 0.2992 0.8650 0.8597 1.0000 1.0000 0.6714 0.6279 0.8384 0.7035 0.7026 0.5662 0.7386
ET-Bert (pkt) 0.9845 0.9824 0.9289 0.9254 0.9872 0.9887 1.0000 1.0000 1.0000 1.0000 0.9704 0.9480 0.9320 0.9181 0.9721
Lens  (flow) 0.9822 0.9808 0.9189 0.9143 0.9867 0.9816 1.0000 1.0000 0.9506 0.9507 0.9712 0.9285 0.9054 0.9061 0.9663
Lens  (pkt) 0.9832 0.9806 0.9349 0.9266 0.9896 0.9880 0.9982 0.9981 0.9885 0.9885 0.9841 0.9658 0.9412 0.9363 0.9784
@@
First, we evaluate the performance of Lens on a variety of traffic understanding tasks. We compare it with the following baselines: FS-Net [21], BiLSTM_Att [43], Datanet [39], DeepPacket [23], YaTC [45], TSCRNN [18], and ET-BERT [19]. In this experiment, we assess different methods with a focus on two levels of data granularity: flow level and packet level. Tables 4 and 5 illustrate the comparison results of our Lens and the baselines in traffic understanding tasks. From the evaluation results, we can conclude with three main insights: (1) Lens  consistently achieves SOTA performance in most tasks, indicating its superiority over all the baselines. (2) With limited testing samples, Lens using packet-level input reaches higher accuracy and F1 than that of using flow-level input. In particular, ET-BERT demonstrates a significant average accuracy difference of 30.13% between these two data granularities. In contrast, our Lens still achieves comparable performance, with only 1.41% variance across two levels of data granularity. (3) Even if without using the CrossPlatform dataset in pre-training, Lens outperforms the baselines in three out of four tasks, especially for tasks 8, 10, and 11, demonstrating its strong generalization ability.
@@
In addition, we conduct experiments to evaluate the performance of Lens on traffic generation tasks by comparing it with the SOTA method, NetShare [44]. In this experiment, two evaluation metrics: Jensen-Shannon Divergence (JSD) and Total Variation Distance (TVD), are used to measure the results. In terms of metrics JSD and TVD, a lower value means better performance, indicating a smaller distribution difference between the generated and ground-truth data. The generation results related to the port number and packet length are illustrated in Tables 6, while the results of IP generation are introduced in Appendix A.1. We draw some conclusions from this table as follows: (1) Lens  outperforms NetShare in most tasks in terms of TVD and JSD, demonstrating its good generation performance. Compared to NetShare, Lens  can generate more high-quality synthetic traffic data by reducing JSD and TVD by 62% and 77% on average, respectively. (2) Lens  consistently outperforms the NetShare in the task of packet length generation under both metrics. (3) Even if our model does not pre-train on the CrossPlatform dataset, it still surpasses the NetShare across most generation tasks in this dataset. More visual results are presented in Appendix A.2.